# Ultralight Dark Matter Constraints from NanoHertz Gravitational Waves

### v1

@@ Abstract
We investigate the impact of ultralight dark matter (ULDM) on the mergers of supermassive black holes (SMBH) and the resulting stochastic gravitational wave background. ULDM is based on exceptionally light particles and yields galactic halos with dense central solitons. This increases the drag experienced by binary SMBH, decreasing merger times and potentially suppressing gravitational radiation from the binary at low frequencies. We develop semi-analytic models for the decay of SMBH binaries in ULDM halos and use current pulsar timing array (PTA) measurements to constrain the ULDM particle mass and its fractional contribution to the dark matter content of the universe. We find a median ULDM particle mass of $7.\times 10^{-22}$ eV and show that scaling relations suggest that the drag remains effective at relatively low ULDM fractions, which are consistent with all other constraints on the model. Consequently, future pulsar timing measurements will be a sensitive probe of any ULDM contribution to the overall dark matter content of the universe.
@@ I Introduction
This paper builds on previous studies of the impact of UDLM on the gravitational wave background from SMBH binaries. In particular, Aghaie et al. [27] impose constraints on the soliton and particle mass under the assumption that the dark matter consists entirely of ULDM and that the SMBHs are larger than the solitons. We generate a synthetic spectrum and also allow for mixed ULDM and CDM models, but their findings are consistent with ours. Separately, two of us (ST and CG) looked at the related issue of drag on SMBH binaries in self-interacting dark matter [28], and this analysis builds on tools developed for that work.
We find that for a particle of around $10^{-21}$eV the drag is sufficient to solve the final parsec problem and to reduce the long wavelength power in the stochastic gravitational wave background, such that the 68% credible interval of $\log_{10}(m/{10^{-22}\mathrm{eV}})=0.84^{+0.44}_{-0.48}$. Interestingly, the drag remains effective if the mass of the central soliton is reduced, as would happen with a mixture of ULDM and regular CDM. As a result, our constraints are largely independent of the assumed ULDM fraction.
@@ II.2 Orbital Decay Rate
where subscript $j$ indicates the other black hole. This provides the torque,
$\frac{\mathrm{d}L_{i}}{\mathrm{d}t}$ $=\frac{\mathrm{d}r_{i}}{\mathrm{d}t}\frac{\mathrm{d}L_{i}}{\mathrm{d}r_{i}}$
$=\frac{\mathrm{d}r_{i}}{\mathrm{d}t}\frac{M_{i}\sqrt{G}}{2}\times\left(\sqrt{\frac{M_{\textrm{enc}_{i}}+\frac{M_{j}^{3}}{\left(M_{i}+M_{j}\right)^{2}}}{r_{i}}}+\right.\left.\frac{\mathrm{d}M_{\textrm{enc}_{i}}}{\mathrm{d}r_{i}}\sqrt{\frac{r_{i}}{M_{\textrm{enc}_{i}}+\frac{M_{j}^{3}}{\left(M_{i}+{M_{j}}\right)^{2}}}}\right)$
$=\frac{\mathrm{d}r_{i}}{\mathrm{d}t}\frac{M_{i}}{2}\left(v_{i}+\frac{4\pi r_{i}^{2}\rho_{i}G}{v_{i}}\right),$ (14)
where in the final line we have made use of spherical symmetry to find ${\mathrm{d}M_{\textrm{enc}_{i}}}/{\mathrm{d}r_{i}}$. This can be equated to the torque due to dynamical friction, $F_{i}r_{i}$, to find ${\mathrm{d}r_{i}}/{\mathrm{d}t}$,
@@
The density in equation (15) depends on the soliton profile, which is “pinched” by the pair of SMBH. We estimate the modified profile by assuming that the total mass of the binary is located at the soliton centre. We additionally assume time-independence (in the radial direction) with the ansatz $\psi(\bm{r},t)=e^{i\beta t}F(r)$ and $\Phi_{\textrm{U}}(\bm{r},t)=\phi(r)$ and the resulting Schrödinger-Poisson equation is
$F^{\prime\prime}(r)+\frac{2}{r}F^{\prime}(r)-2\tilde{\phi}(r)F(r)+2\frac{M_{BH}}{r}F(r)=0,$ (16)
$\tilde{\phi}^{\prime\prime}(r)-4\pi F^{2}(r)+\frac{2}{r}\tilde{\phi}^{\prime}(r)=0,$ (17)
where $\tilde{\phi}(r)=\phi(r)+\beta$ and $F^{\prime}(r)={\mathrm{d}F}/{\mathrm{d}r}$. The equation is solved via the shooting method, with $F(0)=1$, $F^{\prime}(0)=\tilde{\phi^{\prime}}(0)=0$, and $F(r_{\textrm{max}})={\phi}(r_{\textrm{max}})=0$, where $r_{\textrm{max}}$ is a cut-off radius where the soliton profile is negligible. The mass of the soliton can then be found by integrating the profile, and this in turn gives the mass ratio between the black hole and the soliton.
@@ III Gravitational strain spectrum
$h_{c}^{2}(f_{\mathrm{gw}})=\int dM$ $\,dq\,dz\frac{\partial^{3}\eta}{\partial M\partial q\partial z}\cdot\left(-\frac{3}{2}\frac{D}{\dot{D}}\right)$ (27)
$\times 4\pi c\,(1+z)\,d_{c}^{2}$
$\times\frac{32}{5c^{8}}\frac{(G\mathcal{M})^{10/3}}{d_{c}^{2}}(2\pi f_{\mathrm{orb}})^{4/3}.$
For the full model $\dot{D}$ is given by equation (15) and by equation (19) for the simple model – given the ${\mathrm{d}r_{i}}/{\mathrm{d}t}$ we add them to get $\dot{D}$.
Binary separation $D$ is related to the observed gravitational wave frequency $f_{\mathrm{gw}}$, total binary mass $M$, and redshift $z$ through equations (36) and (22). For each combination of ($M,q,z$) we evolve our binary system in 300 steps using $\dot{D}$, starting from a separation of $D=1\mathrm{pc}$, down to $D=\mathrm{ISCO}$, where the ISCO is the innermost stable circular orbit, given by ${6GM}/{c^{2}}$. At each step of the evolution we have a specific value of $D$ which gives us the corresponding $f_{\mathrm{orb}}$, $f_{\mathrm{gw}}$, and the source strain $h_{s}^{2}(f_{\mathrm{orb}})$. This allows us to numerically evaluate the integral in equation (24) for each PTA frequency bin. Following Agazie2023, we set the integration limits to $M:(10^{4}\mathrm{M}_{\odot},10^{12}\mathrm{M}_{\odot})$, $q:(0.001,1)$, and $z:(0.001,10)$. We also require that contributions to the total strain are from binaries that merge within the age of the universe, ensuring that we only consider systems for which fthe inal parsec problem is actually solved in our model.
@@
where $M_{\star 1}$ is the stellar mass of the primary merging galaxy and $q_{\star}$ is the stellar mass ratio of the secondary to the primary galaxy. The relation between $M_{\star 1}$ and $M_{1}$ can be evaluated from equations (3) and (4), with we set $\alpha_{\mu}$ to 1.1. Using these relations $\partial M_{\star 1}/\partial M$ and $\partial q_{\star 1}/\partial q$ are [28]
$\frac{\partial M_{\star 1}}{\partial M}$ $=\frac{1}{\alpha_{\mu}}\frac{M_{\star 1}}{M}$ (29)
$\frac{\partial q_{\star}}{\partial q}$ $=\frac{1}{\alpha_{\mu}}q_{\star}^{1-\alpha_{\mu}}.$
@@
$\log_{10}\left(\Psi_{0}/\mathrm{Mpc}^{-3}\right)=\psi_{0}+\psi_{z}\cdot z^{\prime},$ (32)
$\log_{10}\left(M_{\psi}/M_{\odot}\right)=m_{\psi,0}+m_{\psi,z}\cdot z^{\prime},$
$\alpha_{\psi}=1+\alpha_{\psi,0}+\alpha_{\psi,z}\cdot z^{\prime}.$
@@
The last term in the RHS of equation (30) is $\partial t/\partial z^{\prime}$, can be calculated using the redshift-time relation
@@
While $\partial^{3}\eta/\partial M\partial q\partial z$ can be calculated using the pre-merger redshift $z^{\prime}$ we need the post-merger redshift $z$ for the other terms in the integral in equation (27). This is found from the pre-merger redshift $z^{\prime}$ and the galaxy merger time $T_{\text{gal-gal}}$ by numerically solving for $z_{\mathrm{post-merger}}$ via
@@ IV Parameter Estimation
Our ULDM models, both simple and realistic, have six free parameters. The ULDM specific parameters are the particle mass $m$, which we express as $m/10^{-22}\mathrm{eV}\equiv m_{22}$ and the ULDM fraction, defined in Section II.2. The other parameters are contained within the galaxy stellar mass function (GSMF) $\Psi(M_{\star},z^{\prime})$; $\psi_{0}$ and $m_{\psi,0}$ from equation (32), and the relation between $M_{\rm BH}$ and $M_{\rm bulge}$ (MMB), expressed in equation (3); $\mu$ and $\epsilon_{\mu}$, following Agazie2023. The priors are listed in Table 1.
We generate a library of gravitational strain spectra for SMBH binary mergers occurring within ULDM halos with different parameter combinations. We created 2000 unique combinations of the 6 model parameters following their prior distributions and ranges using Latin hypercube sampling. We generated 2000 realizations of strain spectra for each of the 2000 parameter combinations, as detailed in Section V-C of Ref. [28] and Section 3.2.2 of Agazie2023.
@@ V Results and Analysis
The posterior distributions for the “simple” and “realistic” models are presented in Figures 2 and 3. The 95% region for $\psi_{0}$ matches well with the results of Ref. [46] who found a best fit of $\psi_{0}\sim-2.5$. Our results for ($\psi_{0}$, $m_{\psi,0}$, $\mu$, $\epsilon_{\mu}$) are in agreement with the results from those in Figure 10 of Agazie2023, a phenomenological model with astrophysical priors and with Ref. [28] for the self-interacting dark matter model.
For the realistic model, the median value and 68% credible interval of the ULDM particle mass $m$ is $6.98^{+12.2}_{-4.7}\times 10^{-22}$ eV. For the simple model the allowed range is $1.43^{+4.0}_{-1.2}\times 10^{-22}$ eV. Figure 4 presents the 68% and 95% confidence parameter regions of our models, along with constraints on the ULDM parameter space [29], and we see that these overlap with significant regions that are permitted by observations. For both models, the ULDM particle mass is well constrained, but the fraction is not, confirming the expectations of Section II.4.
@@
Figure 6: Strain spectra for a range of particle masses in both the simple (left- dashed lines) and realistic (right- solid lines) models. Blue lines are the best-fit spectra. Dark green lines correspond to the spectra with best-fit $\mathrm{log}_{10}\left(\frac{m}{10^{-22}eV}\right)-0.2$ and the light green to best-fit $\mathrm{log}_{10}\left(\frac{m}{10^{-22}eV}\right)+0.2$.
@@
Finally, we check the significance of the “signal” of ULDM in the PTA data via Bayesian “Leave one out cross-validation” (LOO-CV) model comparison [48, 49, 50]. This approach is computationally efficient as it needs only the already computed MCMC chains [49]. We compare the simple and realistic ULDM models to the default model, where only gravitational waves contribute to energy loss.
Model name ELPD difference DSE ELPD diff./DSE
GW only 0 0 -
ULDM simple 0.14 3.00 0.05
ULDM realistic 1.91 3.14 0.61
Table 3: LOO-CV model comparison
The “ELPD difference” represents the difference in expected log predictive density between each model and the best model, and “DSE” represents the standard error of the ELPD difference. If the ratio of ELPD difference is greater than 3 standard errors (or DSE), the models are considered to be significantly different. We use the Python package ArviZ [51] to find the results in Table 3. In our case, these ratios are smaller than unity, so there is no meaningful “detection” even though the posteriors on the ULDM parameters are nontrivial. As noted in the Introduction, this is not surprising – the result is driven by the lowest frequency bin, and the statistical power of the deficit itself is not overwhelming.
@@ VI Discussion and Conclusion
We have confirmed that ULDM can contribute significantly to the orbital decay of SMBH binaries at the centres of the largest galaxies. The additional friction remains significant as binaries enter the pulsar timing band, resolving any “final parsec problem” and also suppressing the low-frequency component of the stochastic background, improving the fit to current data. Moreover, this shows that future pulsar timing measurements of the stochastic gravitational wave background will significantly constrain any ULDM contribution to the dark matter component of the universe.
@@
This work is based on an improved semi-analytic model for the dynamical friction on an SMBH binary in a ULDM soliton. This includes a heuristic rescaling of the soliton mass, accounting for both uncertainty in the core-halo relation and scenarios with a fractional ULDM dark matter contribution. Interestingly, scalings within the model yield the prediction that the dynamical friction from ULDM decreases slowly with the ULDM fraction. As a result, the contribution to binary decay remains significant even in scenarios with a relatively small fraction of ULDM.
@@
Because we work directly with a galaxy merger history, we can identify the systems which make the largest contribution to the stochastic background; these are the halos in which the SMBH pair has roughly twice the mass of the soliton. This contrasts with previous work on ULDM-SMBH interactions by Aghaie et al [27], which only considers systems where the black hole mass is at least five times larger than the soliton mass, but our results are consistent with and complementary to theirs.
Constraints on astrophysical parameters in our analyses are in excellent agreement with previous results [34, 28, 46]. Significant pieces of the ULDM parameter space are ruled out by a variety of astrophysical constraints. However, much of the region inside the joint posterior distribution for the ULDM parameters is consistent with all currently available constraints [29].
@@
We developed two models of the dynamical friction - the “realistic” scenario accounts for SMBH binary pinching the soliton profile, whereas the “simple” one does not. The best-fit spectra for both models match the data with a similar level of quality. In both cases, scaling relations in the model mean that drag drops as a small, fractional power of the ULDM fraction. As a result, we cannot put a clear lower bound on the ULDM fraction. That said, the model relies on a number of hypotheses, and the soliton mass is set via the “standard” core-halo relation. Critically, this relation is an almost entirely unexplored relationship in models with a central SMBH, baryons and a mixture of ULDM and CDM. Each of these questions is an important avenue for future investigation. We have also assumed that there is a single ULDM component, which need not be the case, given that the arguments used to produce light scalar fields will typically be consistent with a spectrum with several such species [52].
@@
Constraints on the stochastic background are expected to improve substantially as higher-quality PTA data becomes available in the next few years. If the data continues to suggest an absence of power at low frequencies (relative to the naïve Hellings-Downs curve), it is likely that PTA observations will put tight constraints on the mass of any ULDM particle. Conversely, if this feature is not present in future datasets, PTA observations will put stringent constraints on the overall ULDM parameter space.

### v3

@@ Abstract
We investigate the impact of ultralight dark matter (ULDM) on the mergers of supermassive black holes (SMBH) and the resulting stochastic gravitational wave background. ULDM is based on exceptionally light particles and yields galactic halos with dense central solitons. This increases the drag experienced by binary SMBH, decreasing merger times and potentially suppressing gravitational radiation from the binary at low frequencies. We develop semi-analytic models for the decay of SMBH binaries in ULDM halos and use current pulsar timing array (PTA) measurements to constrain the ULDM particle mass and its fractional contribution to the dark matter content of the universe. We find a median ULDM particle mass of $7\times 10^{-22}$ eV and show that scaling relations suggest that the drag remains effective at relatively low ULDM fractions, which are consistent with all other constraints on the model. Consequently, future pulsar timing measurements will be a sensitive probe of any ULDM contribution to the overall dark matter content of the universe.
@@ I Introduction
This paper builds on previous studies of the impact of UDLM on the gravitational wave background from SMBH binaries. In particular, Aghaie et al. [4] impose constraints on the soliton and particle mass under the assumption that the dark matter consists entirely of ULDM and that the SMBHs are larger than the solitons. We generate a synthetic spectrum and allow for mixed ULDM and CDM models, but their findings are consistent with ours. Tomaselli [50] conducted an analysis of pulsar timing residuals and similarly found a possible suppression at low frequencies due to ULDM drag, further motivating this work. Sarkar [40] used dynamical friction from self-interacting ULDM to impose constraints on the interaction strength, providing a complement to this study. Separately, two of us (ST and CG) looked at the related issue of drag on SMBH binaries in self-interacting dark matter [49], and this analysis builds on tools developed for that work.
We find that for a particle of around $10^{-21}$eV the drag is sufficient to solve the final parsec problem and to reduce the long wavelength power in the stochastic gravitational wave background, such that the 68% credible interval of $\log_{10}(m/{10^{-22}\mathrm{eV}})=0.84^{+0.44}_{-0.48}$. Interestingly, the drag remains effective if the mass of the central soliton is reduced, as would happen with a mixture of ULDM and regular CDM. As a result, our constraints are largely independent of the assumed ULDM fraction, provided a soliton forms.
@@ II.2 Orbital Decay Rate
where subscript $j$ indicates the other black hole. This provides the torque, {align} dLidt=dridtdLidri
= dridt MiG2 ×(Menci+Mj3(Mi+ Mj)2ri+ dMencidririMenci+Mj3(Mi+ Mj)2)
=dridtMi2(v_i+4πri2ρiGvi), where in the final line we have made use of spherical symmetry to find ${\mathrm{d}M_{\textrm{enc}_{i}}}/{\mathrm{d}r_{i}}$. This can be equated to the torque due to dynamical friction, $F_{i}r_{i}$, to find ${\mathrm{d}r_{i}}/{\mathrm{d}t}$,
@@
The density in equation (14) depends on the soliton profile, which is “pinched” by the pair of SMBH. We estimate the modified profile by assuming that the total mass of the binary is located at the soliton centre. We additionally assume time-independence (in the radial direction) with the ansatz $\psi(\bm{r},t)=e^{i\beta t}F(r)$ and $\Phi_{\textrm{U}}(\bm{r},t)=\phi(r)$ and the resulting Schrödinger-Poisson equation is {align} F”(r)+2rF’(r)-2~ϕ(r)F(r)+2MBHrF(r)=0,
~ϕ”(r)-4πF^2(r)+2r~ϕ’(r)=0, where $\tilde{\phi}(r)=\phi(r)+\beta$ and $F^{\prime}(r)={\mathrm{d}F}/{\mathrm{d}r}$. The equation is solved via the shooting method, with $F(0)=1$, $F^{\prime}(0)=\tilde{\phi^{\prime}}(0)=0$, and $F(r_{\textrm{max}})={\phi}(r_{\textrm{max}})=0$, where $r_{\textrm{max}}$ is a cut-off radius where the soliton profile is negligible. The mass of the soliton can then be found by integrating the profile, and this in turn gives the mass ratio between the black hole and the soliton.
@@ III Gravitational strain spectrum
$aligned h_{c}^{2}(f_{\mathrm{gw}})=\int dM&\,dq\,dz\frac{\partial^{3}\eta}{\partial M\partial q\partial z}\cdot\left(-\frac{3}{2}\frac{D}{\dot{D}}\right)\\ &\times 4\pi c\,(1+z)\,d_{c}^{2}\\ &\times\frac{32}{5c^{8}}\frac{(G\mathcal{M})^{10/3}}{d_{c}^{2}}(2\pi f_{\mathrm{orb}})^{4/3}.$ (24)
[…]
@@
[…]
$aligned\frac{\partial M_{\star 1}}{\partial M}&=\frac{1}{\alpha_{\mu}}\frac{M_{\star 1}}{M}\\ \frac{\partial q_{\star}}{\partial q}&=\frac{1}{\alpha_{\mu}}q_{\star}^{1-\alpha_{\mu}}.$ (26)
@@
[…]
@@
[…]
@@
[…]
@@ IV Parameter Estimation
[…]
We generate a library of gravitational strain spectra for SMBH binary mergers occurring within ULDM halos with different parameter combinations. We created 3000 unique combinations of the 6 model parameters following their prior distributions and ranges using Latin hypercube sampling. We generated 2000 realizations of strain spectra for each of the 2000 parameter combinations, as detailed in Section V-C of Ref. [49] and Section 3.2.2 of Agazie2023.
@@ V Results and Analysis
The posterior distributions for the “simple” and “realistic” models are presented in Figures 2 and 3. The 95% region for $\psi_{0}$ matches well with the results of Ref. [5] who found a best fit of $\psi_{0}\sim-2.5$. Our results for ($\psi_{0}$, $m_{\psi,0}$, $\mu$, $\epsilon_{\mu}$) are mainly prior driven and are in agreement with the results from those in Figure 10 of Agazie2023, a phenomenological model with astrophysical priors and with Ref. [49] for the self-interacting dark matter model.
For the realistic model, the median and 68% credible interval of the ULDM particle mass $m$ is $6.53^{+6.7}_{-4.6}\times 10^{-22}$ eV. For the simple model the allowed range is $1.78^{+3.0}_{-1.3}\times 10^{-22}$ eV. Figure 4 presents the 68% and 95% confidence parameter regions of our models, along with constraints on the ULDM parameter space [29], and we see that these overlap with significant regions that are permitted by observations. For both models, the ULDM particle mass is well constrained, but the fraction is not, confirming the expectations of Section II.4.
@@
Figure 6: Strain spectra for a range of particle masses in both the simple (left- dashed lines) and realistic (right- solid lines) models. Blue lines are the best-fit spectra. Dark green lines correspond to the spectra with best-fit $\mathrm{log}_{10}\left(\frac{m}{10^{-22}eV}\right)-0.25$ and the light green to best-fit $\mathrm{log}_{10}\left(\frac{m}{10^{-22}eV}\right)+0.25$.
@@ VI Discussion and Conclusion
We have confirmed that ULDM can contribute significantly to the orbital decay of SMBH binaries at the centers of the largest galaxies. The additional friction remains significant as binaries enter the pulsar timing band, resolving any “final parsec problem” while suppressing the low-frequency component of the stochastic background, improving the fit to current data. A significant fraction of the ULDM parameter space is ruled out by a variety of astrophysical constraints [29]. However, much of the region inside the joint posterior distribution for the ULDM parameters is consistent with all currently available constraints. Moreover, this shows that future pulsar timing measurements of the stochastic gravitational wave background will significantly constrain any ULDM contribution to the dark matter component of the universe.
@@
This work is based on an improved semi-analytic model for the dynamical friction on an SMBH binary in a ULDM soliton. This includes a heuristic rescaling of the soliton mass, accounting for both uncertainty in the core-halo relation and scenarios with a fractional ULDM dark matter contribution. Interestingly, scalings within the model imply that the dynamical friction from ULDM decreases slowly with the ULDM fraction. As a result, the contribution to binary decay remains significant even in scenarios with a relatively small quantity of ULDM.
@@
Because we work directly with a galaxy merger history we can identify the systems which make the largest contribution to the stochastic background; these are the halos in which the SMBH pair has roughly twice the mass of the soliton. This contrasts with previous work on ULDM-SMBH interactions by Aghaie et al. [4], which only considers systems where the black hole mass is at least five times larger than the soliton mass but our results are consistent with and complementary to theirs. Moreover, constraints on astrophysical parameters from our analyses are in excellent agreement with previous results [2, 49, 5], which serves as a check on our model.
@@
We developed two models of the dynamical friction - the “realistic” scenario accounts for the SMBH binary pinching the soliton profile, whereas the “simple” one does not. The best-fit spectra for both models match the data with a similar level of quality. In both cases, scaling relations in the model mean that drag drops as a small, fractional power of the ULDM fraction. As a result, we cannot put a clear lower bound on the ULDM fraction. That said, the model relies on a number of hypotheses, and the soliton mass is set via the “standard” core-halo relation. Critically, this relation is an almost entirely unexplored relationship in models with a central SMBH, baryons and a mixture of ULDM and CDM. Each of these questions is an important avenue for future investigation. We have also assumed that there is a single ULDM component, which need not be the case, given that the arguments used to produce light scalar fields will typically be consistent with a spectrum with several such species [20].
@@
Constraints on the stochastic background are expected to improve substantially as higher-quality PTA data becomes available in the next few years. If the data continues to suggest an absence of power at low frequencies (relative to the simple gravitational-wave–only SMBHB background prediction) PTA observations will identify a preferred region of the ULDM parameter space. Conversely, if this feature is not present in future datasets PTA observations will put stringent constraints on any possible ULDM contribution to the universe.
In the relativistic regime, the possibility of the binary-ULDM system forming a so-called ’gravitational atom’ for certain ULDM mass parameters has been considered [8], wherein excitation of the bound ULDM states produces an additional friction effect that may further impact the gravitational wave signal [22]. This additional effect may change our parameter constraints if added and is an interesting avenue for future work. Conversely, Ref. [14] considers the contribution of three-body scattering and eccentric orbits to the decay of SMBH binaries in the pulsar timing band and it will be critical to check that all environment effects associated with conventional components of the universe are fully accounted for.